OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. a. 1 b. 3 c. 5 d. 6

OpenStudy (anonymous):

80p = 2pr (r+h) 80p/(2p*5)= (5+h) 8=5+h h=3

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

welcome =)

OpenStudy (anonymous):

√96 ^ √32 * √3 ^ √8 √4 ^ ^ √2 * √2 √4 * √2 ^ √2 * √2

OpenStudy (anonymous):

wrong question sorries